OpenStudy (anonymous):

rational expression equations : (3/2k^2) - (1/2k)=(k+2)/(k^2)

OpenStudy (anonymous):

(6-k)/(2k^2) = [2(k+2)]/(2k^2) compare the numerator, you get 6-k = 2k +4 and you can get you k easily

OpenStudy (anonymous):

sorry, i mislook

OpenStudy (anonymous):

it is (3-k) instead of 6-k

OpenStudy (anonymous):

and what do i do next do i mulipy the first fractions of the problem or the (3-k)/(2k^2)=2(k+2)/(2k^2)